OpenStudy (anonymous):

I GIVE MEDALS. You accidentally drop a book down a stairwell. Assuming no air resistance, how fast will the book be moving after 1.1 seconds? A. 10.8 m/s B. 4.9 m/s C. 1.1 m/s D. 9.8 m/s

OpenStudy (anonymous):

A 10.8m/s Since the acceleration of the boook is g(9.8m/s^2), velocity as a function of time is given by v=g*t v=9.8*1.1 v=10.78--->10.8